# Generating Exceptional Knots and Links with Arbitrary Braiding Topology

### 04 Mar 2026 v1

@@ Appendix C Hamiltonian for tight-binding models
For the Trefoil knot $3_{1}$ in the Rolfsen’s table, Torus knot $T(2,3)$], we impose $f_{1}+if_{2}=u^{2}+v^{3}$, then we obtain
@@
$\begin{split}f_{1}&=(\frac{3}{2}P^{2}+Q^{2}-\frac{1}{2})+\frac{P^{2}}{2}cos(2k_{\alpha})+\frac{1}{2}cos(2k_{z})-2PQcos(k_{\alpha})+P^{2}cos(k_{\alpha}\pm k_{\beta})\\ &\quad-\frac{1}{4}sin(3k_{x})+\frac{3}{4}sin(k_{x}\pm 2k_{y})-\frac{3}{4}sin(k_{x}),\\ f_{2}&=\frac{1}{4}sin(3k_{y})+\frac{3}{4}sin(k_{y})+Psin(2k_{z})-2Qsin(k_{z})-\frac{3}{4}[sin(2k_{x}+k_{y})-sin(2k_{x}-k_{y})]\\ &\quad+P[sin(k_{x}+k_{z})-sin(k_{x}-k_{z})+sin(k_{y}+k_{z})-sin(k_{y}-k_{z})],\end{split}$ (C1)
where $k_{\alpha}$ means sum over $\alpha=x,y,z$, and $k_{\alpha}\pm k_{\beta}$ means sum over $\alpha\beta=xy,yz,zx$.
@@
For the Hopf Link $2_{1}^{2}$ in the Rolfsen’s table of link and $L2a1$ in the Thistlethwaite’s link table, Torus knot $T(2,2)$], we impose $f_{1}+if_{2}=u^{2}+v^{2}$, then we obtain
@@
$\begin{split}f_{1}&=(\frac{3}{2}P^{2}+Q^{2}-\frac{1}{2})+\frac{P^{2}}{2}cos(2k_{\alpha})+\frac{1}{2}[-cos(2k_{x})+cos(2k_{y})+cos(2k_{z})]-2PQcos(k_{\alpha})+P^{2}cos(k_{\alpha}\pm k_{\beta}),\\ f_{2}&=Psin(2k_{z})-2Qsin(k_{z})+P[sin(k_{x}+k_{z})-sin(k_{x}-k_{z})+sin(k_{y}+k_{z})-sin(k_{y}-k_{z})]\\ &\quad+[cos(k_{x}-k_{y})-cos(k_{x}+k_{y})].\end{split}$ (C2)
@@
For the Figure-eight knot $4_{1}$ in the Rolfsen’s table, Lemniscate knot $L(2,3,2)$], we substitute Eq. (17) into Eq. (8), then we obtain
$\begin{split}f_{1}&=(96P+36a^{2}Q-64P^{3}-288QP^{2})+16P^{3}cos(3k_{\alpha})-96QP^{2}cos(2k_{\alpha})+[14a^{3}cos(2k_{x})-14a^{3}cos(2k_{y})\\ &\quad-96Qcos(2k_{z})]+(240P^{3}+192Q^{2}P-36a^{2}P-96P)cos(k_{\alpha})+48Pcos(k_{\alpha}\pm 2k_{z})+48P^{3}[cos(k_{\alpha}\pm 2k_{\beta})\\ &\quad+cos(2k_{\alpha}\pm k_{\beta})]-192QP^{2}cos(k_{\alpha}\pm k_{\beta})+96P^{3}cos(k_{x}\pm k_{y}\pm k_{z})+24a^{2}[sin(k_{x}+k_{y}+k_{z})\\ &\quad+sin(k_{x}-k_{y}-k_{z})-sin(k_{x}+k_{y}-k_{z})-sin(k_{x}-k_{y}+k_{z})],\\ f_{2}&=16sin(3k_{z})+(288P^{2}+192Q^{2}-36a^{2}-48)sin(k_{z})+48P^{2}[sin(2k_{\alpha}+k_{z})-sin(2k_{\alpha}-k_{z})]\\ &\quad+96P^{2}[sin(k_{x}+2k_{z})-sin(k_{x}-2k_{z})+sin(k_{y}+2k_{z})-sin(k_{y}-2k_{z})]+96P^{2}[sin(k_{x}\pm k_{y}+k_{z})\\ &\quad-sin(k_{x}\pm k_{y}-k_{z})]-192PQ[sin(k_{\alpha}+k_{z})-sin(k_{\alpha}-k_{z})]+a^{3}[cos(k_{x}-3k_{y})-cos(k_{x}+3k_{y})\\ &\quad+cos(3k_{x}+k_{y})-cos(3k_{x}-k_{y})]+24a^{2}P[cos(k_{x}-2k_{y})-cos(k_{x}+2k_{y})+cos(2k_{x}-k_{y})-cos(2k_{x}+k_{y})]\\ &\quad+24a^{2}P[cos(k_{x}-k_{y}\pm k_{z})-cos(k_{x}+k_{y}\pm k_{z})]-48a^{2}Q[cos(k_{x}-k_{y})-cos(k_{x}+k_{y})].\end{split}$ (C3)
@@
The corresponding tight-binding model is then obtained by inserting $f_{1}$ and $f_{2}$ into Eqs. (1-3). The constructive approach is general and can be similarly applied to derive tight-binding models for other complex knotted and linked ELs. The degree of the semiholomorphic function $F_{a}(u,v,\bar{v})$, defined as the maximum sum of the exponents of $u$, $v$ and $\bar{v}$ in each term of the polynomial, dictates the maximum range of hopping in the tight-binding model. For example, for the Trefoil knot with $F_{a}=u^{2}+v^{3}$, the degree is $3$, leading to the third-nearest-neighbor couplings. Similarly, for the Hopf link with with $F_{a}=u^{2}+v^{2}$, only leading to next-nearest-neighbor couplings. Thus, the longer hopping in the non-Hermitian systems can induce more intricate and complex morphology of ELs.

### 11 Mar 2026 v2

@@ Appendix C Hamiltonian for tight-binding models
For the Trefoil knot $3_{1}$ in the Rolfsen’s table, Torus knot $T(2,3)$], we impose $f_{1}+if_{2}=u^{2}+v^{3}$ and obtain
@@
$\begin{split}f_{1}&=(\frac{3}{2}P^{2}+Q^{2}-\frac{1}{2})+\frac{P^{2}}{2}cos(2k_{\alpha})+\frac{1}{2}cos(2k_{z})-2PQcos(k_{\alpha})+P^{2}cos(k_{\alpha}\pm k_{\beta})\\ &\quad-\frac{1}{4}sin(3k_{x})+\frac{3}{4}sin(k_{x}\pm 2k_{y})-\frac{3}{4}sin(k_{x}),\\ f_{2}&=\frac{1}{4}sin(3k_{y})+\frac{3}{4}sin(k_{y})-2Qsin(k_{z})-\frac{3}{4}[sin(2k_{x}+k_{y})-sin(2k_{x}-k_{y})]\\ &\quad+P\left[sin(k_{\alpha}+k_{z})-sin(k_{\alpha}-k_{z})\right],\end{split}$ (C1)
adopting the summation convention where $k_{\alpha}$ denotes a summation over $\alpha=x,y,z$, and $k_{\alpha}\pm k_{\beta}$ denotes a summation over the index pairs $\alpha\beta=xy,yz,zx$.
@@
For the Hopf Link $2_{1}^{2}$ in the Rolfsen’s table of link and $L2a1$ in the Thistlethwaite’s link table, Torus knot $T(2,2)$], we impose $f_{1}+if_{2}=u^{2}+v^{2}$ and obtain
@@
$\begin{split}f_{1}&=(\frac{3}{2}P^{2}+Q^{2}-\frac{1}{2})+\frac{P^{2}+1}{2}cos(2k_{\alpha})-cos(2k_{x})-2PQcos(k_{\alpha})+P^{2}cos(k_{\alpha}\pm k_{\beta}),\\ f_{2}&=P\left[sin(k_{\alpha}+k_{z})-sin(k_{\alpha}-k_{z})\right]-2Qsin(k_{z})+[cos(k_{x}-k_{y})-cos(k_{x}+k_{y})].\end{split}$ (C2)
@@
For the Figure-eight knot $4_{1}$ in the Rolfsen’s table, Lemniscate knot $L(2,3,2)$], we substitute Eq. (17) into Eq. (8) to obtain
$\begin{split}f_{1}&=(96P+36a^{2}Q-64P^{3}-288QP^{2})+16P^{3}cos(3k_{\alpha})-96QP^{2}cos(2k_{\alpha})+[14a^{3}cos(2k_{x})-14a^{3}cos(2k_{y})\\ &\quad-96Qcos(2k_{z})]+(240P^{3}+192Q^{2}P-36a^{2}P-96P)cos(k_{\alpha})+48Pcos(k_{\alpha}\pm 2k_{z})+48P^{3}[cos(k_{\alpha}\pm 2k_{\beta})\\ &\quad+cos(2k_{\alpha}\pm k_{\beta})]-192QP^{2}cos(k_{\alpha}\pm k_{\beta})+96P^{3}cos(k_{x}\pm k_{y}\pm k_{z})+24a^{2}[sin(k_{x}+k_{y}+k_{z})\\ &\quad+sin(k_{x}-k_{y}-k_{z})-sin(k_{x}+k_{y}-k_{z})-sin(k_{x}-k_{y}+k_{z})],\\ f_{2}&=16sin(3k_{z})+(288P^{2}+192Q^{2}-36a^{2}-48)sin(k_{z})+48P^{2}[sin(2k_{\alpha}+k_{z})-sin(2k_{\alpha}-k_{z})]\\ &\quad+96P^{2}[sin(k_{x}+2k_{z})-sin(k_{x}-2k_{z})+sin(k_{y}+2k_{z})-sin(k_{y}-2k_{z})]+96P^{2}[sin(k_{x}\pm k_{y}+k_{z})\\ &\quad-sin(k_{x}\pm k_{y}-k_{z})]-192PQ[sin(k_{\alpha}+k_{z})-sin(k_{\alpha}-k_{z})]+a^{3}[cos(k_{x}-3k_{y})-cos(k_{x}+3k_{y})\\ &\quad+cos(3k_{x}+k_{y})-cos(3k_{x}-k_{y})]+24a^{2}P[cos(k_{x}-2k_{y})-cos(k_{x}+2k_{y})+cos(2k_{x}-k_{y})\\ &\quad-cos(2k_{x}+k_{y})]+24a^{2}P[cos(k_{x}-k_{y}\pm k_{z})-cos(k_{x}+k_{y}\pm k_{z})]-48a^{2}Q[cos(k_{x}-k_{y})-cos(k_{x}+k_{y})].\end{split}$ (C3)
@@
The corresponding tight-binding models are derived by inserting these expressions for $f_{1}$ and $f_{2}$ into Eqs. (1-3). The constructive approach is general and can be applied similarly to derive tight-binding models for other complex knotted and linked ELs. The maximum range of hopping in the model is determined by the degree of the semiholomorphic polynomial $F_{a}(u,v,\bar{v})$, defined as the greatest sum of exponents of $u$, $v$ and $\bar{v}$ in any single term. For instance, the Trefoil knot with $F_{a}=u^{2}+v^{3}$ (degree $3$) results in couplings up to the third-nearest neighbors, while the Hopf link with with $F_{a}=u^{2}+v^{2}$ (degree $2$) involves only next-nearest-neighbor couplings. Thus, longer-range hopping in non-Hermitian systems can induce more intricate morphologies of ELs.